# Darboux’s theorem in $p$-adic symplectic geometry

### v1

@@ Abstract.
Let $p$ be a prime number. We derive an analog of Moser’s Path Method for $p$-adic analytic manifolds and use it to prove a $p$-adic analog of Darboux’s Theorem. Hence, in the $p$-adic category there are also no local symplectic invariants other than the dimension, as it happens in the real case: locally near any point, a $2n$-dimensional $p$-adic analytic symplectic manifold is symplectomorphic to $((\mathbb{Q}_{p})^{2n},\sum_{i=1}^{n}\mathrm{d}x_{1}\wedge\mathrm{d}y_{i})$. We also prove a $p$-adic version of Weinstein’s generalization of Darboux’s Theorem to neighborhoods of compact manifolds (Darboux-Weinstein’s Theorem). Finally we apply explicitly Darboux’s Theorem to the physical models by Ablowitz-Ladik and Salerno of the Discrete Nonlinear Schrödinger equation.
@@ 1. Introduction
Let $p$ be a prime number. Rigid analytic geometry or $p$-adic geometry is, as described by Lurie [19], roughly speaking a version of the theory of complex manifolds where we replace the complex numbers by the $p$-adic numbers (see also Scholze [31]). Along the same lines, $p$-adic symplectic geometry is grosso modo the study of symplectic manifolds endowed with an atlas and symplectic form that are analytic and have coefficients in the $p$-adic numbers instead of the real numbers.
(Real) symplectic geometry originates in the study of planetary motions in the seventeenth and eighteenth centuries and has deep connections with mathematical physics, analysis/PDEs, representation theory and many other subjects (see [14, 24, 25, 29, 36] for surveys which show the breadth of symplectic geometry). One of the truly fundamental results of the subject is Darboux’s Theorem [10] (1882), whose modern proof uses the so called Moser’s Path Method [22] (1965). Our goal in this paper is to extend Moser’s Path Method to $p$-adic analytic manifolds and use it to prove $p$-adic analogs of Darboux’s Theorem and Darboux-Weinstein’s Theorem, as well as to derive applications to PDEs.
@@
Moser’s Path Method [22], or Moser’s Trick, is a far-reaching tool used in differential geometry which allows us to find diffeomorphisms between differential forms, provided a vector field satisfying some conditions is found. It can be used to prove Darboux’s Theorem [10], which states that any symplectic form on a $2n$-dimensional manifold is locally symplectomorphic to
$\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i},$
where $(x_{1},y_{1},\ldots,x_{n},y_{n})$ are the standard coordinates on $\mathbb{R}^{2n}$ centered at the origin. Darboux’s result can be generalized from neigborhoods of points to neighborhoods of compact submanifolds, this being the content of the Darboux-Weinstein’s Theorem, proved by Alan Weinstein [35, Theorem 4.1]. The precise statement of Moser’s Path Method is as follows.
Let $\ell$ and $k$ be integers such that $0\leqslant k\leqslant\ell$ and $\ell\geqslant 1$. Let $M$ be an $\ell$-dimensional smooth manifold, $m\in M$ and $\{\omega_{t}\}_{t\in[0,1]}$ a family of smooth $k$-forms on $M$. Let $\mathrm{L}$ denote the Lie derivative. Suppose that there exists a smooth time-dependent vector field $\{X_{t}\}_{t\in[0,1]}$ on $M$ such that $\frac{\mathrm{d}}{\mathrm{d}t}\omega_{t}+\mathrm{L}_{X_{t}}\omega_{t}=0$ and $X_{t}(m)=0$ for every $t\in[0,1]$. Then by [22, Theorem 2] there exists an open neighborhood $U$ of $m$ and a smooth family of $p$-adic analytic diffeomorphisms $\{\psi_{t}:U\to\psi_{t}(U)\subset M\}_{t\in[0,1]}$ such that $\psi_{t}^{*}\omega_{t}=\omega_{0}$ for every $t\in[0,1]$.
The condition that $X_{t}(m)=0$ for every $t\in[0,1]$ can be replaced by $M$ being compact, and in this case one obtains a smooth family of diffeomorphisms $\{\psi_{t}:M\to M\}_{t\in[0,1]}$. Compactness is delicate to use in the $p$-adic category and for the purpose of the present paper we will work with the first condition ($X_{t}(m)=0$ for every $t\in[0,1]$).
This means that $\psi_{t}^{*}\omega_{t}$ is constant, and for $t=0$ is $\omega_{0}$, hence it is always $\omega_{0}$. The trick is very useful in differential geometry and it is used for example in [21, page 121], and is also applicable in more general situations even when $M$ is not compact, see [9, Section 2].
Next we extend Moser’s Path Method to the case where the real field is replaced by the field of the $p$-adic numbers $\mathbb{Q}_{p}$. In this new context, Moser’s Path Method cannot be directly applied, because a crucial part of it is the fact that, if the derivative of a smooth function is zero, then the function is constant. In the $p$-adic case this does not hold at all: there are $p$-adic smooth functions from $\mathbb{Q}_{p}$ to $\mathbb{Q}_{p}$, such as
$f(x)=\sum_{n=\operatorname{ord}(x)}^{\infty}a_{n}p^{2n}$
where
$x=\sum_{n=\operatorname{ord}(x)}^{\infty}a_{n}p^{n}$
is the $p$-adic expansion of $x$, with zero derivative everywhere and which are injective.
Definition 1.1 ($p$-adic analytic family).
$T\times A$ $\to B$
$(t,a)$ $\mapsto f_{t}(a)$
is a $p$-adic analytic map.
@@
Then there exists an open subset $U\subset M$ and a $p$-adic analytic family of $p$-adic analytic diffeomorphisms $\{\psi_{t}:U\to\psi_{t}(U)\subset M\}_{t\in\mathbb{Z}_{p}}$, as in Definition 1.1, such that $\psi_{t}^{*}\omega_{t}=\omega_{0}|_{U}$.
Remark 1.2.
Note that the family of $k$-forms $\omega_{t}$ and the vector field $X_{t}$, seen as power series in the variable $t$, must converge in $p^{-d}\mathbb{Z}_{p}$ in order to apply Moser’s Path Method, while the solution $\psi_{t}$ is only obtained for $t\in\mathbb{Z}_{p}$.
1.2. Applications of Moser’s Path Method to symplectic geometry
Theorem A can be used, as happens in the real case, to prove Darboux theorem in the $p$-adic context:
@@ 1.2. Applications of Moser’s Path Method to symplectic geometry
Let $n$ be a positive integer. Let $p$ be a prime number. Let $(M,\omega)$ be a $2n$-dimensional $p$-adic analytic symplectic manifold and $m\in M$. There exist local $p$-adic analytic coordinates $(x_{1},y_{1},\ldots,x_{n},y_{n})$ around $m$ such that
$\omega=\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}.$
It follows from Theorem B that the only local invariant of $p$-adic analytic symplectic manifolds is the dimension. See Figure 1 for an illustration of Theorem B.
@@
Hence, at least locally, real and $p$-adic symplectic geometry are analogous, which stands in strong contrast with other recent discoveries in this new area such as the fact that Gromov’s Nonsqueezing Theorem [15] does not hold in general for $p$-adic symplectic manifolds [8, Theorem B] (see Table 1). Despite the fact that the $p$-adic Darboux’s Theorem has the same statement as its real counterpart, the proof is significantly more involved due to the special nature of the $p$-adic numbers and the behavior of functions defined on them. Using the $p$-adic Moser’s Path Method we can generalize Theorem B to neighborhoods of compact submanifolds, and obtain a $p$-adic analog of the Darboux-Weinstein’s Theorem [35, Theorem 4.1].
symplectic geometry $p$-adic symplectic geometry
Darboux’s Theorem
Darboux-Weinstein’s Theorem
Gromov’s linear nonsqueezing
Gromov’s nonsqueezing ×
Equivariant Gromov’s nonsqueezing
Few normal forms for integrable systems in dimension 44
(7 forms)
Interesting physical models
@@
Let $n$ be a positive integer. Let $p$ be a prime number. Let $M$ be a $2n$-dimensional $p$-adic analytic manifold and let $Q$ be a compact submanifold of $M$. Let $\omega_{0}$ and $\omega_{1}$ be two $p$-adic analytic symplectic forms on $M$ which coincide on $Q$. Then there exist open neighborhoods $U_{0},U_{1}$ of $Q$ and a $p$-adic analytic diffeomorphism $\psi:U_{0}\to U_{1}$ such that $\psi^{*}\omega_{1}=\omega_{0}$.
1.3. Application of the $p$-adic Moser’s Path Method
@@
In the $p$-adic case, we are able to recover the result from [4], which tells us that this symplectic form can be taken into the standard one:
$\omega_{0}=\mathrm{i}\sum_{j\in J}^{n}\mathrm{d}\psi_{j}\wedge\mathrm{d}\bar{\psi}_{j},$
using as a transformation the flow of the vector field
@@
1.4. References in $p$-adic/symplectic geometry and context of the paper
For an introduction to symplectic geometry, we recommend the textbooks and surveys [5, 16, 20, 21, 23, 24], and for its connection to the theory of quantum states see [11, 32, 34]. The role of $p$-adic numbers in physics has been extensively studied, see for example [2, 12, 13, 28, 33]. A construction of the $p$-adic symplectic and Heisenberg groups and the Maslov index can be found in [17, 37]. The present paper is part of a larger program proposed by Voevodsky, Warren and the second author in [26] to develop a $p$-adic version of symplectic geometry with the goal of later implementing it using a proof assistant (in the language of homotopy type theory and Voevodsky’s Univalent Foundations [1, 27]). Within this program the results of our paper are probably the most foundational of all: they clarify the local theory completely. In our previous works in $p$-adic symplectic geometry we focused on matrices/integrable systems [6, 7] and flexibility/rigidity [8].
@@ 1.4. References in $p$-adic/symplectic geometry and context of the paper
Section 2 proves some intermediate results which we need in order to derive Moser’s Path Method and prove Darboux’s Theorem. Section 3 derive the $p$-adic Moser’s Path Method (Theorem A). Section 4 proves the $p$-adic Darboux’s Theorem (Theorem B). Section 5 proves the $p$-adic Darboux-Weinstein’s Theorem (Theorem C). Finally, Section 6 applies Darboux’s Theorem to the $p$-adic version of the Ablowitz-Ladik and Salerno models. The appendix recalls the concepts we need about $p$-adic numbers, $p$-adic analytic functions, $p$-adic analytic manifolds and $p$-adic symplectic geometry.
@@ Acknowledgments
The first author is funded by grant PID2022-137283NB-C21 of MCIN/AEI/ 10.13039/501100011033 / FEDER, UE and by project CLaPPo (21.SI03.64658) of Universidad de Cantabria and Banco Santander. The second author is funded by a FBBVA (Bank Bilbao Vizcaya Argentaria Foundation) Grant for Scientific Research Projects with title From Integrability to Randomness in Symplectic and Quantum Geometry.
The second author thanks the Dean of the School of Mathematical Sciences Antonio Brú and the Chair of the Department of Algebra, Geometry and Topology at the Complutense University of Madrid, Rutwig Campoamor, for their support and excellent resources he is being provided with to carry out the FBBVA project. The second author also thanks Luis Crespo and Francisco Santos for the hospitality during July and August 2025 at the University of Cantabria when this paper was completed.
@@ 2. Technical results on $p$-adic initial value problems, power series and forms on manifolds
In this section we prove four technical lemmas which we will need to prove the main theorems (Theorems A, B, C). We refer to Appendix A for a quick review of the concepts and results from $p$-adic analysis and $p$-adic geometry we need in this section (in particular see Figures 5 and 6 for representations of $p$-adic balls like the ones which appear in Lemma 2.1 below).
@@
Lemma 2.1 is similar in spirit to our previous result [6, Proposition A.9], but that result was weaker and in particular it only concerned one variable.
@@
The following lemma is a $p$-adic version of the Tubular Neighborhood Theorem [3, page 66].
@@
Since every $p$-adic analytic manifold can be written as a disjoint union of domains of charts in the atlas by [6, Corollary 3.2], all of which are $p$-adic balls, the problem reduces to the case when $M$ is a $p$-adic ball. We may assume that this ball is $(\mathbb{Z}_{p})^{\ell}$.
@@
This is well defined because $f$ is a $p$-adic analytic diffeomorphism, and by definition is a $p$-adic analytic family of maps as in Definition 1.1. Note that $\psi_{0}(x)\in Q$ and $\psi_{1}(x)=x$. Now we have, by equations (A.3) and (A.1), that

### v3

@@ Abstract.
Let $p$ be a prime number. We derive an analog of Moser’s Path Method for $p$-adic analytic manifolds and use it to prove a $p$-adic analog of Darboux’s Theorem. Using it as a stepping stone we give a classification of second-countable $p$-adic analytic symplectic manifolds in terms of $p$-adic volume. This is a symplectic version of a classical result of Serre in $p$-adic analytic geometry from 1965. We also prove a $p$-adic version of Weinstein’s generalization of Darboux’s Theorem to neighborhoods of compact manifolds. Finally we find explicitly Darboux’s coordinates for the physical models by Ablowitz-Ladik and Salerno of the Discrete Nonlinear Schrödinger equation.
Symplectic geometry is deeply connected with classical and quantum mechanics, and a main motivation of this paper is to pursue a $p$-adic version of symplectic geometry following recent developments which incorporate the $p$-adic numbers into mathematical and theoretical physics.
@@ 1. Introduction
In this paper we prove $p$-adic analogs of two fundamental results in geometry: Darboux’s local classification of symplectic manifolds (1882) and Serre’s classification of compact $p$-adic analytic manifolds (1965). In fact we will extend Serre’s classification to the noncompact case as well, only assuming that $p$-adic analytic manifolds are second-countable. Our “symplectic” Serre’s classification implies that all second-countable $p$-adic analytic symplectic manifolds admit global standard Darboux coordinates, which stands in very strong contrast with the real case.
We prove these results using a $p$-adic version of the Moser’s Path Method, which we later also apply to study the physical models by Ablowitz-Ladik and Salerno of the Discrete Nonlinear Schrödinger equation.
In fact, symplectic geometry originated in the study of planetary motions and is deeply connected with classical mechanics. A main motivation for the present paper is to pursue a $p$-adic version of sympletic geometry, in the spirit of other important developments in mathematical physics involing the $p$-adic numbers, see Brekke–Freund [2], Dragovich–Khrennikov–Kozyrev–Volovich [14], Dragovich–Khrennikov–Kozyrev–Volovich–Zelenov [15], Vladimirov–Volovich [42], and the references therein. For treatments involving the $p$-adic numbers in string theory, see Brekke–Freund–Olson–Witten [3], Freund–Olson [17], Freund–Witten [18], Fuquen-Tibatá–García-Compeán–Zúñiga-Galindo [19], García-Compeán–López [20], and Volovich [43]. For applications to the theory of black holes, see Chen–Liu–Hung [7]. For a definition of the $p$-adic symplectic group, Heisenberg group and Maslov index, see Hu–Hu [24] and Zelenov [48].
@@
The following is a $p$-adic analytic analog of Moser’s Path Method; see Appendix B for a review of Moser’s Path Method and an explanation of why it cannot be directly applied in the $p$-adic case (the statement below has additional assumptions).
@@
Then there exists an open subset $U\subset M$ and a $p$-adic analytic family of $p$-adic analytic diffeomorphisms $\{\psi_{t}:U\to\psi_{t}(U)\subset M\}_{t\in\mathbb{Z}_{p}}$, such that $\psi_{t}^{*}\omega_{t}=\omega_{0}|_{U}$.
1.2. Applications of $p$-adic Moser’s Path Method to symplectic geometry
1.2.1. Applications to local symplectic geometry
Theorem A can be used, as happens in the real case, to prove Darboux’s Theorem in the $p$-adic context:
@@ 1.2.1. Applications to local symplectic geometry
Let $n$ be a positive integer. Let $p$ be a prime number. Let $(M,\omega)$ be a $2n$-dimensional $p$-adic analytic symplectic manifold and $m\in M$. There exist local $p$-adic analytic coordinates $(x_{1},y_{1},\ldots,x_{n},y_{n})$ around $m$ such that $\omega=\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}.$
@@
It follows from Theorem B that the only local invariant of $p$-adic analytic symplectic manifolds is the dimension. See Figure 1 for an illustration of Theorem B. Using the $p$-adic Moser’s Path Method we can generalize Theorem B to neighborhoods of compact submanifolds, and obtain a $p$-adic analog of the Darboux-Weinstein’s Theorem [45, Theorem 4.1].
@@
Let $n$ be a positive integer. Let $p$ be a prime number. Let $M$ be a $2n$-dimensional $p$-adic analytic manifold and let $Q$ be a compact submanifold of $M$. Let $\omega_{0}$ and $\omega_{1}$ be two $p$-adic analytic symplectic forms on $M$ which coincide on $Q$. Then there exist open neighborhoods $U_{0},U_{1}$ of $Q$ and a $p$-adic analytic diffeomorphism $\psi:U_{0}\to U_{1}$ which fixes $Q$ and such that $\psi^{*}\omega_{1}=\omega_{0}$.
1.2.2. Applications to global symplectic geometry
In what follows, when we say that two $p$-adic manifolds are symplectomorphic, it is assumed that they are so by means of a $p$-adic analytic symplectomorphism.
Theorem D.
Let $v$ be a positive real number. Let $d$ be a positive integer. Then the following statements hold:
  1. (1)
    There exists a noncompact $d$-dimensional $p$-adic analytic manifold $\mathcal{M}_{d,v}\subset(\mathbb{Q}_{p})^{d}$ with $p$-adic volume $v$, defined as follows: if $v=\sum_{k=-\infty}^{m}a_{k}p^{dk}$ is the expression of $v$ in base $p^{d}$ with infinitely many nonzero digits (for example $1$ in base $9$ becomes $0.888\ldots$), then,
    $\mathcal{M}_{d,v}=\bigcup_{k=-\infty}^{m}\bigcup_{i=1}^{a_{k}}\mathrm{B}_{ki}\subset(\mathbb{Q}_{p})^{d}$
    where $\mathrm{B}_{ki}$ is the $d$-dimensional $p$-adic ball defined by
    $\mathrm{B}_{ki}=\mathrm{B}((i_{1}p^{-k-1},i_{2}p^{-k-1},\ldots,i_{d}p^{-k-1}),p^{k}),$
    which are all pairwise disjoint, with $i=\sum_{j=0}^{d-1}i_{j+1}p^{j}$ and $0\leqslant i_{j}\leqslant p-1$. See Figure 5 for the balls $\mathrm{B}_{ki}$.
  2. (2)
    If, in addition, $v$ is rational and its denominator is a power of $p$, then there also exists a compact $d$-dimensional $p$-adic analytic manifold $\mathcal{M}_{d,v}^{\prime}\subset(\mathbb{Q}_{p})^{d}$ with $p$-adic volume $v$, defined as follows: if $v=\sum_{k=\ell}^{m}a_{k}p^{dk}$ is the expression of $v$ in base $p^{d}$ with finitely many nonzero digits, which exists because $v$ is rational and its denominator is a power of $p$, then
    $\mathcal{M}_{d,v}^{\prime}=\bigcup_{k=\ell}^{m}\bigcup_{i=1}^{a_{k}}\mathrm{B}_{ki}\subset(\mathbb{Q}_{p})^{d}.$
Endow $\mathcal{M}_{d,v}$ and $\mathcal{M}_{d,v}^{\prime}$ with the standard symplectic form $\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}\xi_{i}$. If in the definition of $\mathcal{M}_{d,v}$ and $\mathcal{M}_{d,v}^{\prime}$ we change the balls by other pairwise disjoint balls of the same radius the resulting $p$-adic analytic manifolds are symplectomorphic.
Now we turn to the symplectic version of Serre’s result Theorem C.1. The $p$-adic volume which we use here is the $p$-adic measure with respect to the form $\omega^{n}$, where $\omega$ is the symplectic form, analogously to the definition in the real case.
Theorem E (Symplectic Serre’s classification of $p$-adic analytic manifolds).
Let $n$ be a positive integer. Let $p$ be a prime number. Let $(M,\omega)$ be a second-countable $2n$-dimensional $p$-adic analytic symplectic manifold. Let $v$ be the volume of $(M,\omega)$ and let $\mathcal{M}_{2n,v}$ and $\mathcal{M}_{2n,v}^{\prime}$ be as defined in Theorem D. Endow $(\mathbb{Q}_{p})^{2n}$, $\mathcal{M}_{2n,v}$ and $\mathcal{M}_{2n,v}^{\prime}$ with the standard symplectic form $\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}\xi_{i}$.
  1. (1a)
    If $M$ is noncompact and $v=\infty$, $(M,\omega)$ is symplectomorphic to $(\mathbb{Q}_{p})^{2n}$.
  2. (1b)
    If $M$ is noncompact and $v$ is finite, $(M,\omega)$ is symplectomorphic to $\mathcal{M}_{2n,v}$.
  3. (1c)
    If $M$ is compact, then $v$ is a rational number with a power of $p$ as denominator and $(M,\omega)$ is symplectomorphic to $\mathcal{M}_{2n,v}^{\prime}$.
  4. (2)
    $\mathcal{M}_{2n,v}$ and $\mathcal{M}_{2n,v^{\prime}}$ are symplectomorphic if and only if $v=v^{\prime}$. The same happens with $\mathcal{M}_{2n,v}^{\prime}$ and $\mathcal{M}_{2n,v^{\prime}}^{\prime}$.
Recall that $\mathcal{M}_{2n,v}$ and $\mathcal{M}_{2n,v}^{\prime}$ are defined as unions of balls, hence Theorem E is a symplectic version of Serre’s Classification (Theorem C.1); the difference is that the symplectic case needs an unbounded number of balls. Theorem E implies that there exists a $p$-adic analytic symplectic embedding $\varphi:(M,\omega)\to((\mathbb{Q}_{p})^{2n},\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}\xi_{i}).$ Hence, there always exist global Darboux coordinates.
Theorem F.
Let $n$ be a positive integer. Let $p$ be a prime number. Let $(M,\omega)$ and $(M^{\prime},\omega^{\prime})$ be second-countable $2n$-dimensional $p$-adic analytic symplectic manifolds with the same $p$-adic volume, finite or infinite, and which are both compact or both noncompact. Then $(M,\omega)$ is symplectomorphic to $(M^{\prime},\omega^{\prime})$.
Theorem F stands in very strong contrast with real symplectic geometry where there is an extensive theory of other invariants, such as symplectic capacities.
1.2.3. Application of the $p$-adic Moser’s Path Method to the Ablowitz-Ladik and Salerno models
@@
In the $p$-adic case, we are able to recover the result from [5], which tells us that this symplectic form can be taken into the standard one $\omega_{\mathrm{std}}=\mathrm{i}\sum_{j\in J}^{n}\mathrm{d}\psi_{j}\wedge\mathrm{d}\bar{\psi}_{j},$ using as a transformation the flow of the vector field
@@
Section 2 proves some intermediate results which we need in order to derive Moser’s Path Method and prove Darboux’s Theorem. Section 3 derives the $p$-adic Moser’s Path Method (Theorem A). Section 4 proves the $p$-adic Darboux’s Theorem (Theorem B). Section 5 proves the $p$-adic Darboux-Weinstein’s Theorem (Theorem C). Section 6 proves Theorems D and E. Finally, Section 7 applies Darboux’s Theorem to the $p$-adic version of the Ablowitz-Ladik and Salerno models. The appendices recall the concepts we need about $p$-adic numbers, $p$-adic analytic functions, $p$-adic analytic manifolds and $p$-adic symplectic geometry (Appendix A), the real Moser’s Path Method (Appendix B) and the classical Serre’s classification (Appendix C).
Funding
The first author is funded by grant PID2022-137283NB-C21 of MCIN/AEI/ 10.13039/501100011033 / FEDER, UE. The second author is funded by a FBBVA (Bank Bilbao Vizcaya Argentaria Foundation) Grant for Scientific Research Projects with title From Integrability to Randomness in Symplectic and Quantum Geometry.
@@ Acknowledgments
The second author thanks the Dean of the School of Mathematical Sciences Antonio Brú and the Chair of the Department of Algebra, Geometry and Topology at the Complutense University of Madrid, Rutwig Campoamor, for their support and excellent resources he is being provided with to carry out the FBBVA project. The second author also thanks Luis Crespo and Francisco Santos for the hospitality during July and August 2025 at the University of Cantabria when part of this paper was written.
@@ 2. Technical results on $p$-adic initial value problems, power series and forms on manifolds
In this section we prove four technical lemmas which we will need to prove the main theorems (Theorems A, B, C, D, E, F). We refer to Appendix A for a quick review of the concepts and results from $p$-adic analysis and $p$-adic geometry we need in this section (in particular see Figures 7 and 8 for representations of $p$-adic balls like the ones which appear in Lemma 2.1 below).
@@
Lemma 2.1 is similar in spirit to our previous result [8, Proposition A.9], but that result was weaker and in particular it only concerned one variable.
@@
The following lemma is a $p$-adic version of the Tubular Neighborhood Theorem [4, page 66].
@@
Since every $p$-adic analytic manifold can be written as a disjoint union of domains of charts in the atlas by [8, Corollary 3.2], all of which are $p$-adic balls, the problem reduces to the case when $M$ is a $p$-adic ball. We may assume that this ball is $(\mathbb{Z}_{p})^{\ell}$.
@@
This is well defined because $f$ is a $p$-adic analytic diffeomorphism, and by definition is a $p$-adic analytic family of maps. Note that $\psi_{0}(x)\in Q$ and $\psi_{1}(x)=x$. Now we have, by equations (A.3) and (A.1), that
@@
The last two results in this section concern symplectomorphisms between unions of balls.
Lemma 2.9.
Let $p$ be a prime number. Let $r$ be a $p$-adic absolute value. The $2n$-dimensional $p$-adic ball of radius $pr$ is a disjoint union of $p^{2n}$ $2n$-dimensional $p$-adic balls of radius $r$.
Proof.
If $r=p^{-k}$ and $m$ is a point in $\mathrm{B}(0,pr)=\mathrm{B}(0,p^{1-k})$, then the digits at the right of the position k1k-1 in all coordinates of $m$ are 0. We can divide the ball in $p^{2n}$ balls according to the digit at position $k$ of each one of the $2n$ coordinates; these balls are a translation of $\mathrm{B}(0,p^{-k})$, as we wanted. ∎
Lemma 2.10.
Let $p$ be a prime number. Let $B,B_{1},\ldots,B_{k}$ be $p$-adic balls such that $\operatorname{vol}(B)=\sum_{i=1}^{k}\operatorname{vol}(B_{i})$, where $\operatorname{vol}$ denotes $p$-adic volume. Then $B$ is symplectomorphic to $\bigcup_{i=1}^{k}B_{i}$.
Proof.
Let $v$ be the smallest volume among the balls $B_{i}$. If $\operatorname{vol}(B)=v$, there is nothing to prove. Otherwise, $\operatorname{vol}(B)$ is a multiple of $p^{2n}v$ and the same happens with all volumes greater than $v$. Hence, there are at least $p^{2n}$ balls with volume $v$. Using Lemma 2.9, we can unify these $p^{2n}$ balls in one with the immediately higher radius. Now we repeat the operation until there is only one ball. ∎